OpenStudy (anonymous):

is each function odd, even or neither? f(x)=1-x^2/2+x^2

OpenStudy (jamesj):

A function is odd if f(-x) = - f(x) And it is even if f(-x) = f(x) So to find out whether your function is odd or even, evaluate f(-x) and see if it is equal to either f(x) or -f(x). What do you find?

OpenStudy (unklerhaukus):

you might want to simply first \[f(x)=1-{x^2\over2}\]

OpenStudy (jamesj):

I think the function is \[ f(x) = \frac{1-x^2}{2 + x^2} \] but Danyel should clarify.

OpenStudy (unklerhaukus):

i diden't even think of that

OpenStudy (anonymous):

this is what I came up with yes